DU PONT U we 6 Determine the payback period. Cash Flow ما دل - 1o
Answer #1

PV of cashflow should turn equal to 100 to achieve payback.

Hence PV= A*(1-(1+r)^-n/r

Or 100=25*(1-(1+6%)^-n)/6%

Or, 100/25*0.06=1-1.06^-n

Or, 1.06^-n=1-0.24

Or, -nlog 1.06=log 0.76

Or, n=-log0.76/log1.06

Or, n=0.1192/0.0253

Or n=4.71

Hence payback is 4.71 years